Prehistoric cave paintings were discovered in a cave in France. The paint contained 35 %35% of the original​ carbon-14. Use the exponential decay model for​ carbon-14, Upper A equals Upper A 0 e Superscript negative 0.000121 tA=A0e−0.000121t​, to estimate the age of the paintings. The paintings are approximately nothing years old. ​(Round to the nearest​ integer.)
Answer #1

The exponential decay model for​ carbon-14 is A = A0e -0.000121t where A0 is the initial quantity of carbon-14and A is the quantity of carbon-14 left after t years.

Here, A = 35 % of A0 = 0.35 A0 so that 0.35 A0 = A0 e -0.000121t or, e -0.000121t = 0.35.

Now, on taking natural log of both the sides, we get ln(e -0.000121t) = ln (0.35) or, -0.000121t = -1.049822125 so that t =1.049822125 / 0.000121 = 8676.215905 = 8676 ( on rounding off to the nearest integer).

Thus, the paintings are approximately 8676 years old.